Circular No. 5672

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bm IN ANONYMOUS GALAXY Jean Mueller reports her discovery of an apparent supernova (blue mag

∼ 18) located 11′′ south of the nucleus of a galaxy at α = 7^{h}29^{m}30^{s}.40, δ = +50^{o}34′51′′.8 (equinox 1950.0). The supernova’s position has end figures 30^{s}.41, 38′′.4. SN 1992bm was found on a plate taken Nov. 25 UT by C. Brewer and Mueller with the 1.2-m Oschin Telescope in the course of the second Palomar Sky Survey. N. Reid observed the object Dec. 1 with the Palomar 1.5-m telescope (+ CCD camera) and reports that the object has very blue colors V = 18.3, V–I = +0.04.

PERIODIC COMET SWIFT-TUTTLE (1992t) J. G. Hills, Los Alamos National Laboratory, notes: “The nongravita-

tional forces observed on this comet and the intensity and long duration of its meteor stream suggest that large amounts of material have been shed from this object. At the comet’s next return, the earth may pass large pieces of comet debris as it approaches the central core of the Perseid meteor stream. At an impact velocity of 60 km/s, calved pieces ∼ 100 m in diameter could cause airblast damage at the earth’s surface. Particularly after perihelion, observers should search for evidence of calved pieces along the orbit of the comet, especially near the point where the earth is predicted to cross the orbit of the comet in 2126 (ΔT ∼ +15 days).”

U. Fink, University of Arizona, writes: “Narrow-band filter imaging and long-slit spectroscopy of P/Swift-Tuttle were obtained during Nov. 24-29 UT by M. Hicks, U. Fink, M. Hoffmann, and N. Chanover with the 1.54-m Catalina Site Telescope (+ Lunar and Planetary Laboratory CCD and spectrometer system) of the University of Arizona Observatories. The comet showed three secondary nuclei (B, C, D), whose rough separation, position angle, and magnitude difference with respect to the primary nucleus (A) are: B, 12′′.4 (11 900 km), 44^{o}, 3.2; C, 19′′.5 (18 600 km), 70^{o}, 5.2; D, 25′′.6 (24 500 km), 42^{o}, 6.2. Over the six days of observation, no motion with respect to the primary nucleus could be discerned. In addition to the secondary nuclei, very strong jet activity was observed. The secondary nuclei showed their own distinct comae and associated jet structure.”

CORRIGENDA

On IAUC 5558, title, for ‘340-G38’ read ‘ESO 340-G38’; on IAUC 5660, ‘Supernova 1992bl in ESO 291-G11’, line 12, for 1992bk read 1992bl

1992 December 8 Daniel W. E. Green